OpenStudy (anonymous):

Which of the following has a slope of –3 and a y-intercept of –4?

OpenStudy (anonymous):

Where are the options? :)

OpenStudy (anonymous):

what are the choices?

OpenStudy (anonymous):

What they said ^^^

OpenStudy (anonymous):

A. y = –3x + 4 B. y = –3x – 4 C. y = 3x + 4 D.y = 3x – 4

OpenStudy (anonymous):

remember in the equation y=mx+b (m) is the slope it determines the rise over run while the y intercept is the point on the y axis :)

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Yes, m is slope or -3 in this case. The y-intercept is the LAST number or b in y - mx + b. BE SURE TO CARRY THE NEGATIVE SIGN.

OpenStudy (anonymous):

@greenlegodude57 ok

OpenStudy (anonymous):

@sana1234 ok

OpenStudy (anonymous):

how do i give the medal out

OpenStudy (anonymous):

yes B

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

CLICK BEST RESPONSE

OpenStudy (anonymous):

Click "Best Response"

OpenStudy (luigi0210):

Please don't give out direct answers..

OpenStudy (anonymous):

WHY? @Luigi0210

OpenStudy (anonymous):

@Luigi0210 We explained it to him. He understands.

OpenStudy (luigi0210):

Yea, you're suppose to give the help without the answer..

OpenStudy (anonymous):

Doesn't matter, next time he can do a problem like this himself.

OpenStudy (anonymous):

O